OpenStudy (anonymous):

What is the Thevenin equivalent resistance of the small-signal circuit, as seen from the output port? Write an algebraic expression for this resistance in terms of the parameters K and VT, and the bias voltage VIN.

OpenStudy (anonymous):

1/(K*(VIN-VT))

OpenStudy (anonymous):

Rth = 1/gm = 1/(K*(VIN-VT))